# Structure, short-range order, and phase stability of the AlxCrFeCoNi high-entropy alloy: Insights from a perturbative, DFT-based analysis

### v1

@@ I Introduction
Our own methodology [36, 37, 38, 39, 40, 41] for modelling the phase stability of these alloys falls into the last of the aforementioned categories, and is based on a perturbative analysis of the internal energy of the disordered alloy as evaluated within the CPA. Notably, compared to some alternative CPA-based schemes, we carefully and fully include the effects of charge rearrangement due to an applied chemical perturbation [36], which turns out to be essential when modelling many HEA systems [37, 38]. A schematic of the workflow is visualised in Fig. 1. In summary, the process begins by performing a self-consistent DFT calculation to evaluate the internal energy of the disordered solid solution. A perturbative analysis then examines the stability of this disordered phase to applied chemical fluctuations within a concentration wave formalism, as pioneered for binary alloys by Khachaturyan [42] and Gyorffy and Stocks [43]. Via application of a Landau-type theory, it is possible to infer chemical orderings directly, obtaining the preferred ordered structure, the temperature at which it emerges, and information about which chemical species preferentially sit on which sublattice. In addition to the results of the Landau theory, from this perturbative analysis it is also possible to recover a representation of the internal energy of the alloy in terms of effective pair interactions. This effective Hamiltonian can then be used in computationally inexpensive atomistic simulations to study the phase behaviour of a system below any initial ordering temperature in detail. Previously, we have made use of the Metropolis Monte Carlo algorithm for this purpose [37, 38, 39, 40, 41], but in the present work we instead apply the nested sampling algorithm [44, 45], a sampling technique which is fully predictive, truly unbiased, and capable of simultaneously exploring competing free energy basins. It is therefore well-suited to studying the complex phase behaviour of high-entropy alloys.
@@ II.1 Underlying Crystal Lattice
Using the all-electron HUTSEPOT code [59], we compute the total energy-per-atom of the disordered AlxCrFeCoNi system in both disordered fcc (A1) and bcc (A2) structures as a function of $x$ in the range $0\leq x\leq 2$, and calculate the difference in energy between the two structures. The results are shown in Fig. 2. Lattice parameters are set to be consistent with their experimental values [17]. For values of $x$ where the lattice type is not observed experimentally (e.g. bcc, $x=0$; fcc, $x=2$) we perform a volume-conserving transformation from fcc to bcc structures (or vice-versa) to obtain a lattice parameter. Full details of these calculations are described in Sec. IV.4. Dashed lines on the figure indicate the regions where, experimentally, the system is observed to form an fcc lattice, a mixed fcc/bcc phase, or purely the bcc lattice [11]. It can be seen that the KKR-CPA calculations successfully capture the transition of the underlying crystal lattice from fcc to bcc with increasing $x$. We deduce that the experimentally observed fcc$+$bcc coexistence region originates from the small difference in energy between the two structures in the region $0.5\leq x\leq 1$, leading to competition between the two structures.
@@ II.2 Electronic Structure
Proceeding, we consider the electronic DoS for both fcc and bcc structures, as shown in Fig. 4. The most obvious contrast between the two lattice types is the difference in the contribution to the total DoS made by the narrow, $3d$ bands associated with the transition metals Cr, Fe, Co, and Ni. In the bcc structure, particularly at lower values of $x$, there is a notable dip in the total and species-resolved DoS curves at intermediate energies. This is associated with splitting of $e_{g}$ (higher energy) and $t_{2g}$ (lower energy) states in this structure. In our calculations, this splitting is increased in the bcc structure compared to the fcc structure. Sample $(l,m)$-resolved DoS plots evidencing this are given in the Supplemental Material [60]. As for the BSF plots of Fig. 3, another key feature is the broad, $sp$ band associated with Al, which can be seen hybridising with the $d$ bands of the transition metals, indicated by localised peaks in its species-resolved DoS. With increasing $x$, a peak in the total DoS at lower energies is maintained in the bcc structure, while in the fcc structure the total DoS appears gradually pushed to higher energies. Another important point to note is the decreasing valence electron concentration (VEC) with increasing Al content. For $x=0$ (i.e. the composition CrFeCoNi) the VEC is 8.25 $e$/atom, while for $x=1$ (i.e. the composition AlCrFeCoNi), the VEC is 7.2 $e$/atom. We interpret the combination of the aforementioned band mechanisms (splitting of $3d$ states, $p$-$d$ hybridisation) and reducing VEC as the electronic origins of the stabilisation of the bcc phase with increasing Al content.
@@
In addition to these features, a comment should also be made about charge transfer between chemical species in the alloyed system. In an alloy, it is typically a good approximation for the average lattice parameter of the system to be determined from a weighted average of the atomic volumes of each element present, in accordance with Vegard’s law [61]. However, this means that charge from the Wigner-Seitz cells of ‘large’ atoms in the alloy typically spills over into the Wigner-Seitz cell of the ‘small’ atoms in the alloy. In these systems, we find that Al, the atom with the largest relative atomic volume, typically ‘loses’ some electrons to atoms such as Co and Ni, which have smaller associated atomic volumes. For example, for bcc AlCrFeCoNi ($x=1$), a site occupied by an Al atom (atomic number 13) has an associated charge surrounding it of 12.84 $e$, while Ni (atomic number 28) has an associated charge of 28.13 $e$. (Full details of all relative charge transfers between elements can be found in the potential files in the repository associated with this study.) In previous work, we have found these relative atomic size differences and associated charge transfers between elements to play a key role in determining atomic ordering tendencies as charge rearranges itself in response to an applied chemical perturbation [38]. It is therefore vital that these effects are fully considered in the perturbative stability analysis [36].
@@
Figure 6: The partially ordered structures for AlxCrFeCoNi predicted by the linear response theory for (a) $x=0$ an L12 ordering, (b) $x=0.5$ a D022 ordering, and (c) $x=1$ a B2 ordering. We note that in the case of (b), the L12 and D022 orderings are very close in energy; D022 ordering is only just favoured over L12 in our calculations.
@@ II.3 Perturbative Stability Analysis
We use the self-consistent KKR-CPA one-electron potentials and associated electron densities as the starting point for a linear response calculation assessing the energetic cost of chemical perturbations to the homogeneous, disordered solid solution. These perturbations are naturally expressed within the concentration wave formalism [42, 43]. In this approach, we write an inhomogeneous set of site-wise concentrations (Eq. 1) as a perturbation to the homogeneous site-wise concentrations (Eq. 2),
@@
where $\Delta c_{\alpha}(\mathbf{k})$ represents a static concentration wave with wavevector $\mathbf{k}$ and chemical polarisation $\Delta c_{\alpha}$. Examples of how these concentration waves can be used to describe various ordered structures imposed on the fcc and bcc lattices are given in Sec. IV.
@@
where we define $S^{(2)}_{i\alpha;j\alpha^{\prime}}$ for notational convenience. Again, owing to the translational symmetry of the underlying lattice, it is convenient to inspect the lattice Fourier transform of this quantity, writing $S^{(2)}_{\alpha\alpha^{\prime}}(\mathbf{k})$. This is then an $s\times s$ matrix which is a function of the wavevector, $\mathbf{k}$, where $s$ is the total number of chemical species in the alloy. Its evaluation involves self-consistently solving a set of coupled equations in terms of various CPA-related quantities, the most computationally demanding part of which is a convolution integral over the irreducible Brillouin zone (IBZ) [36]. Evaluation of this term has been discussed extensively in earlier works, so we omit these details here for brevity [36, 37, 38, 39, 40, 41].
@@
where $V_{i\alpha;j\alpha^{\prime}}$ represents the energy associated with an atom of species $\alpha$ on site $i$ interacting with an atom of species $\alpha^{\prime}$ on site $j$. It can be shown that, within our formalism, the Fourier transforms of $S^{(2)}_{\alpha\alpha^{\prime}}(\mathbf{k})$ represent a best choice of effective pair interaction, $V_{i\alpha;j\alpha^{\prime}}$, for use in this model [36]. So, by applying an inverse Fourier transform to the $S^{(2)}_{\alpha\alpha^{\prime}}(\mathbf{k})$ data, we can recover an effective pair interaction for use in subsequent atomistic simulations.
@@ II.4 Inferred Chemical Orderings
From the reciprocal space $S^{(2)}_{\alpha\alpha^{\prime}}(\mathbf{k})$ data, it is possible to reconstruct the chemical stability matrix, $\Psi^{-1}_{\alpha\alpha^{\prime}}(\mathbf{k})$, which assesses the stability of the system to an applied chemical perturbation [36]. This quantity can be thought of as analogous to the Hessian matrix evaluated at a stationary point of the free energy surface of the system. At high temperatures, this matrix is positive definite, and the homogeneous solid solution is stable to any applied chemical perturbation. However, with decreasing temperature, we expect that, at some temperature $T_{\text{ord}}$ and for some wavevector $\mathbf{k}_{\text{ord}}$, an eigenvalue of this matrix will pass through zero, and a chemical ordering can be inferred. The wavevector defining the ordering, $\mathbf{k}_{\text{ord}}$, suggests the ordered structure which will emerge, while the eigenvector itself, $\Delta c_{\alpha}$, informs us of which chemical species partition themselves onto which sublattice(s).
Tabulated in 1 are inferred ordering temperatures, wavevectors, and concentration wave polarisations for the AlxCrFeCoNi system for a range of values of $x$ for both a bcc and fcc underlying lattice. Some examples of the associated predicted ordered structures are given in Fig. 6. We emphasise that, as these ordering temperatures are computed within a single-site theory, they are expected to marginally overestimate the exact transition temperatures [36].
@@ II.5 Atomistic Simulations
In a system with many constituent elements and potentially competing chemical interactions, the above partial orderings inferred by the Landau-type theory may not convey a full picture of the phase behaviour of the system. To address this, we inverse Fourier transform the reciprocal space $S^{(2)}_{\alpha\alpha^{\prime}}(\mathbf{k})$ data to recover the pairwise, real-space interaction of Eq. 6. (Our effective pair interactions, $V_{i\alpha;j\alpha^{\prime}}$, are provided in the repository associated with this work, and are also tabulated in the Supplementary Material.) We find that an interaction limited to the first six coordination shells of the lattice fits the reciprocal space data with extremely good accuracy. Interactions are largest on the first coordination shell and tail off quickly with increasing distance; there are no particularly long-ranged interactions present in our analysis of these systems, although interactions between Al and other elements present are significant up to and including the third coordination shell of the lattice in both fcc and bcc cases. Our fitted interactions are provided in the database associated with this study, and are both tabulated and visualised in the Supplemental Material [60].
@@
The above results, as well as the results inferred directly via the concentration wave analysis in Sec. II.4, are broadly consistent with other theoretical studies, in addition to experimental works. Ref. [25] used embedded atom model (EAM) potentials and Monte Carlo simulations to equilibrate structures at a temperature of 500 K. These authors found short-range order followed by an L12 ordering driven by Al with increasing Al-concentration when the underlying lattice was fcc, and a strong B2 ordering driven by Al when the underlying lattice was bcc. Ref. [27] found similar SRO tendencies using another EAM potential, finding Al-Fe and Cr-Cr nearest-neighbour pairs favoured, and Al-Al and Al-Cr pairs disfavoured. In Ref. [68], lattice-based Monte Carlo simulations using nearest-neighbour pair interactions were employed, and it was found that, when the underlying lattice was bcc, their simulations decomposed into B2 Al-Ni-Co rich phases and disordered bcc Cr-Fe phases. Similarly, Ref. [70] reports combined experimental and theoretical results, and finds that the system decomposes into B2 Ni-Al and Cr-rich phases when the underlying lattice is bcc. That our results are in alignment with these earlier studies reassures us of the predictive capabilities of our modelling approach.

### v2

@@ I Introduction
Our own methodology [36, 37, 38, 39, 40, 41, 42] for modelling the phase stability of these alloys falls into the last of the aforementioned categories, and is based on a perturbative analysis of the internal energy of the disordered alloy as evaluated within the CPA. Notably, compared to some alternative CPA-based schemes, we carefully and fully include the effects of charge rearrangement due to an applied chemical perturbation [36], which turns out to be essential when modelling many HEA systems [37, 38]. A schematic of the workflow is visualised in Fig. 1. In summary, the process begins by performing a self-consistent DFT calculation to evaluate the internal energy of the disordered solid solution. A perturbative analysis then examines the stability of this disordered phase to applied chemical fluctuations within a concentration wave formalism, as pioneered for binary alloys by Khachaturyan [43] and Gyorffy and Stocks [44]. Via application of a Landau-type theory, it is possible to infer chemical orderings directly, obtaining the preferred ordered structure, the temperature at which it emerges, and information about which chemical species preferentially sit on which sublattice. In addition to the results of the Landau theory, from this perturbative analysis it is also possible to recover a representation of the internal energy of the alloy in terms of effective pair interactions. This effective Hamiltonian can then be used in computationally inexpensive atomistic simulations to study the phase behaviour of a system below any initial ordering temperature in detail. Previously, we have made use of the Metropolis Monte Carlo algorithm for this purpose [37, 38, 39, 40, 41, 42], but in the present work we instead apply the nested sampling algorithm [45, 46], a sampling technique which is fully predictive, truly unbiased, and capable of simultaneously exploring competing free energy basins. It is therefore well-suited to studying the complex phase behaviour of high-entropy alloys.
@@ II.1 Underlying Crystal Lattice
Using the all-electron HUTSEPOT code [60], we compute the total energy-per-atom of the disordered AlxCrFeCoNi system in both disordered fcc (A1) and bcc (A2) structures as a function of $x$ in the range $0\leq x\leq 2$, and calculate the difference in energy between the two structures. The results are shown in Fig. 2. Lattice parameters are set to be consistent with their experimental values [17]. For values of $x$ where the lattice type is not observed experimentally (e.g. bcc, $x=0$; fcc, $x=2$) we perform a volume-conserving transformation from fcc to bcc structures (or vice-versa) to obtain a lattice parameter. Full details of these calculations are described in Sec. IV.4.
Dashed lines on the figure indicate the regions where, experimentally, the system is observed to form an fcc lattice, a mixed fcc/bcc phase, or purely the bcc lattice, where the experimental data is taken from Ref. [11]. These authors report the observed phases for the AlxCrFeCoNi high-entropy alloy in two conditions: as-cast, and homogenised at 1100 C for 24 hrs. In the as-cast samples, it is found that fcc-bcc phase coexistence occurs in the region $0.880.45<x<$ 0.88, while in the homogenised samples the coexistence region is enlarged to $1.170.30<x<$ 1.17. We use the larger of these two windows to plot the dashed lines in Fig. 2.
It can be seen that the KKR-CPA calculations successfully capture the transition of the underlying crystal lattice from fcc to bcc with increasing $x$. We deduce that the experimentally observed fcc$+$bcc coexistence region originates from the small difference in energy between the two structures in the region $0.5\leq x\leq 1$, leading to competition between the two structures.
@@ II.2 Electronic Structure
Proceeding, we consider the electronic DoS for both fcc and bcc structures, as shown in Fig. 4. The most obvious contrast between the two lattice types is the difference in the contribution to the total DoS made by the narrow, $3d$ bands associated with the transition metals Cr, Fe, Co, and Ni. In the bcc structure, particularly at lower values of $x$, there is a notable dip in the total and species-resolved DoS curves at intermediate energies. This is associated with splitting of $e_{g}$ (higher energy) and $t_{2g}$ (lower energy) states in this structure. In our calculations, this splitting is increased in the bcc structure compared to the fcc structure. Sample $(l,m)$-resolved DoS plots evidencing this are given in the Supplemental Material [61]. As for the BSF plots of Fig. 3, another key feature is the broad, $sp$ band associated with Al, which can be seen hybridising with the $d$ bands of the transition metals, indicated by localised peaks in its species-resolved DoS. With increasing $x$, a peak in the total DoS at lower energies is maintained in the bcc structure, while in the fcc structure the total DoS appears gradually pushed to higher energies.
Another important point to note is the decreasing valence electron concentration (VEC) with increasing Al content. For $x=0$ (i.e. the composition CrFeCoNi) the VEC is 8.25 $e$/atom, while for $x=1$ (i.e. the composition AlCrFeCoNi), the VEC is 7.2 $e$/atom. Varying VEC has previously been noted as an important factor in determining the underlying lattice type of high-entropy alloys in both experimental [62, 63] and theoretical and computational [64, 65] studies. For example, Guo et al. [62] found that, experimentally, alloys with a VEC greater than 8 $e$/atom were typically fcc, while those with a VEC less than 6.87 $e$/atom were typically bcc, findings which are broadly in line with ours in the present study. Attempts have also been made to provide theoretical insight into these experimental observations. Where states at the Fermi surface are free-electron-like in character, for example in alloys of noble metals, the VEC determines where the Fermi sphere intersects with the Brillouin zone boundary [66, 64]. As the first Brillouin zones for fcc and bcc lattices are different in shape, it can be argued that the differing shapes of the two Brillouin zones allow for more electrons to be accommodated in lower energy levels for one structure over the other, thus determining the relative stability [66]. However, for systems such as those of the present study, where the $d$-bands are partially filled, such simple arguments do not apply. A more sophisticated argument was made by Poletti and Battezzati [64], who employed a modified Friedel model to account for partial filling of $d$-bands. These authors found that, when the VEC was less than 7.5 $e$/atom, the bcc structure was favoured in their modelling, while when the VEC was greater than 7.5 $e$/atom, it was the fcc structure which was favoured. This is in reasonable agreement with our own findings. However, we stress that for systems such as the AlxCrFeCoNi alloy considered in this work, where magnetism plays an important role and the system’s $d$-bands are heavily smeared by the disorder, it is challenging to identify a single particular electronic mechanism driving the transition from fcc to bcc lattice type with increasing Al concentration.
We interpret the combination of the aforementioned band mechanisms (splitting of $3d$ states, $p$-$d$ hybridisation) and reducing VEC as the electronic origins of the stabilisation of the bcc phase with increasing Al content.
@@
In addition to the above features, a comment should also be made about charge transfer between chemical species in the alloyed system. In an alloy, it is typically a good approximation for the average lattice parameter of the system to be determined from a weighted average of the atomic volumes of each element present, in accordance with Vegard’s law [67]. However, this means that charge from the Wigner-Seitz cells of ‘large’ atoms in the alloy typically spills over into the Wigner-Seitz cell of the ‘small’ atoms in the alloy. In these systems, we find that Al, the atom with the largest relative atomic volume, typically ‘loses’ some electrons to atoms such as Co and Ni, which have smaller associated atomic volumes. For example, for bcc AlCrFeCoNi ($x=1$), a site occupied by an Al atom (atomic number 13) has an associated charge surrounding it of 12.84 $e$, while Ni (atomic number 28) has an associated charge of 28.13 $e$. (Full details of all relative charge transfers between elements can be found in the potential files in the repository associated with this study.) Such charge transfer from Al to the other elements present in the composition is reminiscent of the notion of aluminium having a negative effective valency when alloyed with transition metals, as first proposed by Raynor [66]. Such a notion is supported by the propensity of aluminium to form intermetallic compounds with a range of $3d$ transition metals [68]. In previous work, we have found these relative atomic size differences and associated charge transfers between elements to play a key role in determining atomic ordering tendencies as charge rearranges itself in response to an applied chemical perturbation [38]. It is therefore vital that these effects are fully considered in the perturbative stability analysis [36, 41].
@@
Figure 6: The partially ordered structures for AlxCrFeCoNi predicted by the linear response theory for (a) $x=0$ an L12 ordering, (b) $x=0.5$ a D022 ordering, and (c) $x=1$ a B2 ordering. We note that in the case of (b), the L12 and D022 orderings are very close in energy; D022 ordering is only just favoured over L12 in our calculations. We emphasise that these orderings are only the initial orderings inferred from the solid solution by our linear response theory, and that further chemical orderings and/or eventual segregation can occur at lower temperatures. This aspect is examined in Sec. IV.2.
@@ II.3 Perturbative Stability Analysis
We use the self-consistent KKR-CPA one-electron potentials and associated electron densities as the starting point for a linear response calculation assessing the energetic cost of chemical perturbations to the homogeneous, disordered solid solution. This approach is discussed in detail in Sec. IV.1 and in references  [37, 38, 39, 40, 41, 42], so for brevity we only outline the key details here. Chemical perturbations are naturally expressed within the concentration wave formalism [43, 44]. In this approach, we write an inhomogeneous set of site-wise concentrations (Eq. 1) as a perturbation to the homogeneous site-wise concentrations (Eq. 2),
@@
where $\Delta c_{\alpha}(\mathbf{k})$ represents a static concentration wave with wavevector $\mathbf{k}$ and chemical polarisation $\Delta c_{\alpha}$. Examples of how these concentration waves can be used to describe various ordered structures imposed on the fcc and bcc lattices are given in Sec. IV. The wavevectors associated with a given chemical ordering are, naturally, those associated with the superstructure peaks which would be observed in a scattering experiment.
@@
where we define $S^{(2)}_{i\alpha;j\alpha^{\prime}}$ for notational convenience. Again, owing to the translational symmetry of the underlying lattice, it is convenient to inspect the lattice Fourier transform of this quantity, writing $S^{(2)}_{\alpha\alpha^{\prime}}(\mathbf{k})$. (Relevant definitions of lattice Fourier transforms are provided in Sec. IV.5.) $S^{(2)}_{\alpha\alpha^{\prime}}(\mathbf{k})$ is then an $s\times s$ matrix which is a function of the wavevector, $\mathbf{k}$, where $s$ is the total number of chemical species in the alloy. Its evaluation involves self-consistently solving a set of coupled equations in terms of various CPA-related quantities, the most computationally demanding part of which is a convolution integral over the irreducible Brillouin zone (IBZ) [36]. Evaluation of this term has been discussed extensively in earlier works, so we omit these details here for brevity [36, 37, 38, 39, 40, 41, 42].
@@
where $V_{i\alpha;j\alpha^{\prime}}$ represents the energy associated with an atom of species $\alpha$ on site $i$ interacting with an atom of species $\alpha^{\prime}$ on site $j$. It can be shown that, within our formalism, the Fourier transforms of $S^{(2)}_{\alpha\alpha^{\prime}}(\mathbf{k})$ represent a best choice of effective pair interaction, $V_{i\alpha;j\alpha^{\prime}}$, for use in this model [36]. (Note that, on account of the homogeneity of the disordered solid solution, we assume that the real-space interactions are translationally invariant.) So, by applying an inverse Fourier transform to the $S^{(2)}_{\alpha\alpha^{\prime}}(\mathbf{k})$ data, we can recover an effective pair interaction for use in subsequent atomistic simulations.
@@ II.4 Inferred Chemical Orderings
From the reciprocal space $S^{(2)}_{\alpha\alpha^{\prime}}(\mathbf{k})$ data, it is possible to reconstruct the chemical stability matrix, $\Psi^{-1}_{\alpha\alpha^{\prime}}(\mathbf{k})$, which assesses the stability of the system to an applied chemical perturbation [36]. This quantity can be thought of as analogous to the Hessian matrix evaluated at a stationary point of the free energy surface of the system. At high temperatures, this matrix is positive definite, and the homogeneous solid solution is stable to any applied chemical perturbation. However, with decreasing temperature, we expect that, at some temperature $T_{\text{ord}}$ and for some wavevector $\mathbf{k}_{\text{ord}}$, an eigenvalue of this matrix will pass through zero, and a chemical ordering can be inferred. The wavevector defining the ordering, $\mathbf{k}_{\text{ord}}$, suggests the ordered structure which will emerge, while the eigenvector itself, $\Delta c_{\alpha}$, informs us of which chemical species partition themselves onto which sublattice(s). This aspect is discussed in detail in Sec. IV.1, as well as in earlier works [37, 38].
Tabulated in 1 are inferred ordering temperatures, wavevectors, and concentration wave polarisations for the AlxCrFeCoNi system for a range of values of $x$ for both a bcc and fcc underlying lattice. Some examples of the associated predicted ordered structures are given in Fig. 6. We emphasise that, as these ordering temperatures are computed within a single-site theory, they are expected to marginally overestimate the exact transition temperatures [36]. Further, we stress that these orderings are only the initial orderings inferred from the solid solution by our linear response theory, and that further chemical orderings and/or eventual segregation can occur at lower temperatures. This aspect is examined in Sec. IV.2, where atomistic simulations are performed down to room temperature.
@@ II.5 Atomistic Simulations
In a system with many constituent elements and potentially competing chemical interactions, the above partial orderings inferred by the Landau-type theory may not convey a full picture of the phase behaviour of the system. Indeed, many single-phase high-entropy alloys are expected to eventually become metastable with decreasing temperature. To address this aspect, we inverse Fourier transform the reciprocal space $S^{(2)}_{\alpha\alpha^{\prime}}(\mathbf{k})$ data to recover the pairwise, real-space interaction of Eq. 6. (Our effective pair interactions, $V_{i\alpha;j\alpha^{\prime}}$, are provided in the repository associated with this work, and are also tabulated in the Supplementary Material.) We find that an interaction limited to the first six coordination shells of the lattice fits the reciprocal space data with extremely good accuracy. Interactions are largest on the first coordination shell and tail off quickly with increasing distance; there are no particularly long-ranged interactions present in our analysis of these systems, although interactions between Al and other elements present are significant up to and including the third coordination shell of the lattice in both fcc and bcc cases. Our fitted interactions are provided in the database associated with this study, and are both tabulated and visualised in the Supplemental Material [61].
@@
The above results, as well as the results inferred directly via the concentration wave analysis in Sec. II.4, are broadly consistent with other theoretical studies, in addition to experimental works. Ref. [25] used embedded atom model (EAM) potentials and Monte Carlo simulations to equilibrate structures at a temperature of 500 K. These authors found short-range order followed by an L12 ordering driven by Al with increasing Al-concentration when the underlying lattice was fcc, and a strong B2 ordering driven by Al when the underlying lattice was bcc. Ref. [27] found similar SRO tendencies using another EAM potential, finding Al-Fe and Cr-Cr nearest-neighbour pairs favoured, and Al-Al and Al-Cr pairs disfavoured. In Ref. [75], lattice-based Monte Carlo simulations using nearest-neighbour pair interactions were employed, and it was found that, when the underlying lattice was bcc, their simulations decomposed into B2 Al-Ni-Co rich phases and disordered bcc Cr-Fe phases. Similarly, Ref. [77] reports combined experimental and theoretical results, where experimental samples for a range of Al concentrations are annealed at temperatures between 1250 and 700 C. These authors find that the system decomposes into B2 Ni-Al and Cr-rich phases when the underlying lattice is bcc for temperatures between 900 and 700 C. That our results are in alignment with these earlier studies reassures us of the predictive capabilities of our modelling approach.